# Non-Invasive Suicide Risk Prediction Through Speech Analysis

### v2

@@ I INTRODUCTION
To this end, machine learning (ML) based approaches and their use for the diagnostic assessment of mental illness by medical professionals have been explored, offering a helpful addition in clinical emergency settings. In previous work, primarily electronic health records (EHR) obtained from hospitals, meta-information such as character traits, or text-based data, e. g., from Twitter, were employed. Walsh et al. [walsh2017predicting] explored an EHR approach to predict suicidal risk and analyzed the risk over time from 7 days to 720 days before the attempted suicide. Their experiments showed that the accuracy increases the closer the date gets to the suicide attempt, achieving a best area under the curve (AUC) of $84\,\%$. Another study that utilized longitudinal EHR data for suicide risk prediction determined short- and long-term risk factors, leading to an AUC of up to $86\,\%$ [su2020machine].
@@
In addition, previous studies explored text-based suicide risk prediction. Roy et al. [roy2020machine] trained several neural networks using posts from Twitter data related to psychological conditions such as burden, anxiety, or loneliness. Subsequently, they trained a random forest model for predicting suicidal ideation cases using the outputs of the previously trained networks, achieving an AUC of $88\,\%$. An extensive linguistic analysis was conducted by Homan et al. [homan2022linguistic]. Their findings show that suicidal thoughts contain more intensifiers and superlatives. In contrast, suicidal behaviors utilized more often pronouns, fewer modifiers, and numerals as well as more prepend words.
@@ II DATASET
We have collected a unique speech-based dataset from $20$ subjects undergoing emergency admission to the Department of Psychiatry at the BLINDED FOR REVIEW. These recordings encompass three distinct types of speech activities: (i) reading of three short narratives [^2] The three short stories were (1) “Der Nordwind und Die Sonne” (the story of ‘the north wind and the sun’), “Gleich am Walde” (the story of ‘right next to the forest’), and “Der Hund und das Stück Fleisch” (the story of ‘the dog and the piece of meat’). widely used within phonetics and speech pathology, each story was read twice; (ii) picture description where each patient describes a picture in their words; and (iii) isolated vowels production (/a:/, /e:/, /i:/, /o:/, /u:/). Each subject contributed 8 recordings after admission to the emergency room (referred to as baseline), and some consented to record an additional 8 upon discharge from hospital care. In our experiments, we restrict the analysis to the baseline samples to allow for a consistent validation procedure. The study doctor assessed the suicide risk level of each patient using the Likert scale [1 – 6].
@@ III-A Segmentation
As a first step in the audio-based analysis of suicide risk, we split the speech recordings into smaller segments – sentences for the picture descriptions and neutral texts, and individual vowels for the vowel production task. This process is automated in whisperX [bain2022whisperx] by force-aligning transcriptions generated by Whisper [Radford23-RSR] to the outputs of a wav2vec phoneme recognition model. We then cut the audio samples at the detected sentence or vowel boundaries. Statistics about the resulting audio segments are provided in Table I.
@@ III-B1 Speech and Acoustic Functionals
We utilize openSMILE [Eyben10-OTM] to extract two sets of features – a comprehensive set of 6 37363736\,3736 373 functionals (ComParE) and eGeMAPS, a smaller selection of features (88 dimensional) crafted for speech emotion recognition.
@@ III-B2 Deep Audio Representations
We enhance our feature set by incorporating DeepSpectrum features [Amiriparian17-SSC], proven to be effective across various speech-based health-related and affective computing tasks [gerczuk2023noise, amiriparian2021ai, amiriparian2023muse]. Additionally, we include embeddings from three distinct pretrained audio Transformers: wav2vec-large [baevski2020wav2vec], and two fine-tuned versions, one on German [grosman2021xlsr53-large-german], and the other on the MSP-Podcast [lotfian2017building] specifically fine-tuned for speech emotion recognition [Wagner23-DOT]. We chose these latter two Transformers to explore their effectiveness on German data and to assess whether a model fine-tuned for emotion tasks can capture relevant patterns in the speech of individuals at risk of suicide. For DeepSpectrum features, we first transfer audio signals into Mel-spectrogram plots with $128$ Mels and forward them through a DenseNet201 architecture (weights pre-trained on ImageNet). Afterward, we extract the activations of the penultimate fully connected, resulting in a $1920$ dimensional feature set. For all wav2vec features, we extract $1024$-dimensional embeddings by passing the speech recordings through the pretrained wav2vec models. These embeddings are obtained by averaging the resulting representations from the final layer of the models.
@@ IV RESULTS
TABLE III: Metadata and Fusion Results: The first column shows the results of models trained solely with metadata, while the subsequent columns represent the results of fusing metadata with the best-performing speech model. F2 to F10 is derived by incrementally adding one metadata to the previous one. The best overall result is highlighted with gray shading.
Figure 1: Comparison of models trained solely with metadata (blue) and models with metadata fused with features from different speech recordings: all speech data (green), speech data exclusively from picture descriptions (red), reading neutral texts (purple), and distinct vowels (orange). Feature types F1–F10: cf., caption of Table III
@@ V CONCLUSIONS AND FUTURE WORK
We have introduced a non-invasive, speech-based approach for automatic suicide risk assessment. Leveraging a novel dataset of speech recordings from $20$ patients, we extracted six sets of audio representations, including interpretable speech and acoustic features, DeepSpectrum features, and embeddings from pretrained audio Transformers. We showed that our model can recognize the degree of suicidality (high/low) from speech with a balanced accuracy of $66.2\,\%$, and when the patients’ metadata is fused with our speech model, a balanced accuracy of up to $94.4\,\%$ can be achieved.

### v3

@@ I INTRODUCTION
To this end, machine learning (ML)-based approaches and their use for the diagnostic assessment of mental illness by medical professionals have been explored, offering a helpful addition in clinical emergency settings. In previous work, primarily electronic health records (EHR) obtained from hospitals, meta-information such as character traits, or text-based data, e. g., from Twitter, were employed. Walsh et al. [walsh2017predicting] explored an EHR approach to predict suicidal risk and analyzed the risk over time from 7 days to 720 days before the attempted suicide. Their experiments showed that the accuracy increases the closer the date gets to the suicide attempt, achieving the best area under the curve (AUC) of $84\,\%$. Another study that utilized longitudinal EHR data for suicide risk prediction determined short- and long-term risk factors, leading to an AUC of up to $86\,\%$ [su2020machine].
@@
In addition, previous studies explored text-based suicide risk prediction. Roy et al. [roy2020machine] trained several neural networks using posts from Twitter data related to psychological conditions such as burden, anxiety, or loneliness. Subsequently, they trained a random forest model for predicting suicidal ideation cases using the outputs of the previously trained networks, achieving an AUC of $88\,\%$. An extensive linguistic analysis was conducted by Homan et al. [homan2022linguistic]. Their findings show that suicidal thoughts contain more intensifiers and superlatives. In contrast, suicidal behaviors utilized more often pronouns, fewer modifiers and numerals, as well as more prepend words.
@@ II DATASET
We have collected a unique speech-based dataset from $20$ subjects undergoing emergency admission to the Department of Psychiatry at the District Hospital Augsburg, Germany. These recordings encompass three distinct types of speech activities: (i) reading of three short narratives [^2] The three short stories were (1) “Der Nordwind und Die Sonne” (the story of ‘the north wind and the sun’), “Gleich am Walde” (the story of ‘right next to the forest’), and “Der Hund und das Stück Fleisch” (the story of ‘the dog and the piece of meat’). widely used within phonetics and speech pathology, each story was read twice; (ii) picture description where each patient describes a picture in their words; and (iii) isolated vowels production (/a:/, /e:/, /i:/, /o:/, /u:/). Each subject contributed eight recordings after admission to the emergency room (referred to as baseline), and some consented to record an additional eight upon discharge from hospital care. In our experiments, we restrict the analysis to the baseline samples to allow for a consistent validation procedure. The study doctor assessed the suicide risk level of each patient using the Likert scale [1 – 6].
@@ III-A Segmentation
As a first step in the audio-based analysis of suicide risk, we split the speech recordings into smaller segments – sentences for the picture descriptions and neutral texts and individual vowels for the vowel production task. This process is automated in whisperX [bain2022whisperx] by force-aligning transcriptions generated by Whisper [Radford23-RSR] to the outputs of a wav2vec phoneme recognition model. We then cut the audio samples at the detected sentence or vowel boundaries. Statistics about the resulting audio segments are provided in Table I.
@@ III-B1 Speech and Acoustic Functionals
We utilize openSMILE [Eyben10-OTM] to extract two sets of features – a comprehensive set of 6,373 functionals (ComParE) and eGeMAPS, a smaller selection of features (88 dimensional) crafted for speech emotion recognition.
@@ III-B2 Deep Audio Representations
We enhance our feature set by incorporating DeepSpectrum features [Amiriparian17-SSC], proven to be effective across various speech-based health-related and affective computing tasks [gerczuk2023noise, amiriparian2021ai, amiriparian2023muse]. Additionally, we include embeddings from three distinct pretrained audio Transformers: wav2vec-large [baevski2020wav2vec] and two fine-tuned versions, one on German [grosman2021xlsr53-large-german] and the other on the MSP-Podcast [lotfian2017building] specifically fine-tuned for speech emotion recognition [Wagner23-DOT]. We chose these latter two Transformers to explore their effectiveness on German data and to assess whether a model fine-tuned for emotion tasks can capture relevant patterns in the speech of individuals at risk of suicide. For DeepSpectrum features, we first transfer audio signals into Mel-spectrogram plots with $128$ Mels and forward them through a DenseNet201 architecture (weights pre-trained on ImageNet). Afterward, we extract the activations of the penultimate fully connected, resulting in a $1920$-dimensional feature set. For all wav2vec features, we extract $1024$-dimensional embeddings by passing the speech recordings through the pretrained wav2vec models. These embeddings are obtained by averaging the resulting representations from the final layer of the models.
@@ IV RESULTS
TABLE III: Metadata and Fusion Results: The first column shows results of models trained with only metadata, while subsequent columns show results of fusing metadata with the top speech model. F2 to F10 are generated by incrementally adding metadata. The best result is shaded gray.
Figure 1: Comparison of models trained solely with metadata (blue) and models with metadata fused with features from different speech recordings: all speech data (green), speech data exclusively from picture descriptions (red), reading neutral texts (purple), and distinct vowels (orange). Feature types F1–F10: cf. Table III
@@ V CONCLUSIONS AND FUTURE WORK
We have introduced a non-invasive, speech-based approach for automatic suicide risk assessment. Leveraging a novel dataset of speech recordings from $20$ patients, we extracted six sets of audio representations, including interpretable speech and acoustic features, DeepSpectrum features, and embeddings from pretrained audio Transformers. We showed that our model can recognize the degree of suicidality (high/low) from speech with a balanced accuracy of $66.2\,\%$, and when it is fused with patients’ metadata, a balanced accuracy of up to $94.4\,\%$ can be achieved.